OpenStudy (anonymous):

f(x)=1/2 x +1 a. Find the Inverse f(x) and name it g(x)

OpenStudy (anonymous):

CHANGE F(X) TO 'Y' Y=1/2X + 1 NOW SWITCH THE PLACES OF X AND Y X=1/2Y + 1 NOW, YOUR TURN.. REARRANGE THE EQUATION IN TERMS OF Y

OpenStudy (ccswims):

"inverse" means a perpendicular slope g(x) = 2x + b (we don't know what the y-intercept is so we don't put anything for b) g(x) = 2x

OpenStudy (ccswims):

2 is the inverse of 1/2

OpenStudy (ccswims):

hope this helped :)

OpenStudy (anonymous):

@chibi_robo3 i got y=2x-2

OpenStudy (anonymous):

@Chibi_Robo3

OpenStudy (anonymous):

x=1/2y + 1 x-1=1/2y 2x-2=y now change y=g(x) therefore, g(x)=2x-2 yep seems right :)

OpenStudy (anonymous):

sorry for the late response, i had to do something